OpenStudy (anonymous):

How many hotdogs and how many cans of soda were sold? Mitchell works at a concession stand selling hotdogs and cans of soda. A hotdog costs $3.00 and a can of soda costs $1.00. One night, he notices that he sold a total of 350 items and collected a total of $624.

OpenStudy (imstuck):

First you have to relate the number of hot dogs and cans of soda to the total number of items sold. This will be equation 1. Then you relate the money part of it...how much each hot dog sells for, how much each can of soda sells for, then have those relate to the total amount of money made. This is the second equation. You will have a system of equations for which you will solve for the variables to get your answer. First one relating the number of each sold looks like this:

OpenStudy (imstuck):



OpenStudy (imstuck):

Then we relate the money that each costs to how much was actually made, which is $624. The number of hot dogs, which we don't know, at $3 apiece, is represented by 3h, which means that each hot dog costs $3. The cans of soda sell for $1 apiece, so that can be represented by 1s (or just s). Adding what they cost together is equal to 624. That second equation looks like this:

OpenStudy (imstuck):



OpenStudy (imstuck):

Now you have a system of equations that needs to be solved for both h and s. Do this one variable at a time. I recommend doing it using the substitution method. Are you familiar with it? I'm sure you are if you are doing these in class.

OpenStudy (anonymous):

Yes, I am. So far I've gotten to... hotdogs=h cans of soda=s s + h = 350 3h + 1s = 624 s = 350 - h 3h + 1(350 - h) = 624 3h + 350 - h = 624 2h = 974 h = 487

OpenStudy (anonymous):

But i just wasn't able to find a number for s because i kept getting a negative result

OpenStudy (anonymous):

@IMStuck

OpenStudy (texaschic101):

s + h = 350 --> s = 350 - h 3h + s = 624 3h + 350 - h = 624 -- subtract 350 from both sides 3h - h = 624 - 350 -- combine like terms 2h = 274 h = 274/2 h = 137 s + h = 350 s + 137 = 350 s = 350 - 137 s = 213 check.. 3h + s = 624 3(137) + 213 = 624 411 + 213 = 624 624 = 624 (correct) 137 hotdogs (h) and 213 sodas (s) were sold

OpenStudy (texaschic101):

what you did was added the 350 instead of subtracting it